AVAC is the only pharmaceutical firm producing a Vaccine. 


The Demand Curve for its product is Qd = 250 – 50 P         

where P is Price and Q are packs of vaccines in ‘000


                        Total Cost Function estimated by the firm is     TC = 15 + 0.5Q      

where Q is monthly output.


a.     What is the market structure of AVAC? State its characteristics.   


b.     To maximize profit,           


(i)             What will be the optimum price and how many packs of Vaccine should the firm produce and sell per month?

(ii)           If this number of packs is produced and sold, what will be the firm’s monthly profit?                                                                     


c.     Using available information, draw  AVAC’s demand, marginal revenue and marginal cost curves in a graph and clearly label thefirm’s profit maximizing price, quantity and profit. Do you observe any welfare loss? If so, also indicate and label the area on the graph.


d.     Assume all other pharmaceutical firms in the market start producing the Vaccine and the market becomes competitive. What will be the impact on price and marginal revenue?Would the market structure of the firm remain the same? Support your answer with help of the firm’s graph.



1
Expert's answer
2021-01-31T18:59:29-0500

a) Since AVAC is the only pharmaceutical firm producing a Vaccine, the market structure of AVAC is monopoly. The characteristics of monopoly include profit maximizer, price maker, high barriers to entry, single seller, and price discrimination.

b)

(i) Let's write the inverse demand function:


"P=5-0.02Q."

Let's write the total and marginal revenue:


"TR=PQ=(5-0.02Q)Q,""MR=5-0.04Q."

Let's find the price and quantity that maximizes profits:


"MR=MC=\dfrac{dTC}{dQ},""5-0.04Q=0.5,""Q=112.5"

Substituting "Q" into the inverse demand function, we get the price:


"P=\$2.75"

(ii) Finally, let's calculate the the firm’s monthly profit:


"\pi=TR-TC,""\pi=(5-0.02\cdot112.5)\cdot112.5-15-0.5\cdot112.5=\$238.12"

(c) Let's draw AVAC’s demand, marginal revenue and marginal cost curves in a graph:



Here, D (blue curve) is the demand, MR (orange curve) is the marginal revenue, MC (grey curve) is the marginal cost. The quantity and price maximizing the firms profit also labeled. The welfare loss (which is exist when the monopoly presented on market) depicted as orange shaded area.

d) In case of competitive market P=MC:


"5-0.02Q=0.5,""Q=225."

Then, substituting "Q" in the inverse demand function, we get:


"P=\$0.5"

Then, the firm's market structure becomes competitive, in this case there is no welfare loss. Let's draw AVAC’s demand, marginal revenue and marginal cost curves in a graph: